### v2

@@ Abstract
Recently, the scale of transformers has grown rapidly, which introduces considerable challenges in terms of training overhead and inference efficiency in the scope of task adaptation. Existing works, namely Parameter-Efficient Fine-Tuning (PEFT) and model compression, have separately investigated the challenges. However, PEFT cannot guarantee the inference efficiency of the original backbone, especially for large-scale models. Model compression requires significant training costs for structure searching and re-training. Consequently, a simple combination of them cannot guarantee accomplishing both training efficiency and inference efficiency with minimal costs. In this paper, we propose a novel Parallel Yielding Re-Activation (PYRA) method for such a challenge of training-inference efficient task adaptation. PYRA first utilizes parallel yielding adaptive weights to comprehensively perceive the data distribution in downstream tasks. A re-activation strategy for token modulation is then applied for tokens to be merged, leading to calibrated token features. Extensive experiments demonstrate that PYRA outperforms all competing methods under both low compression rate and high compression rate, demonstrating its effectiveness and superiority in maintaining both training efficiency and inference efficiency for large-scale foundation models. Our code will be released to the public.
@@ 1 Introduction
Vision transformers [15] have made a profound impact across various domains of computer vision, such as image classification [15, 6, 56, 65, 52, 14], object detection [4, 71, 40, 43, 61, 16], image segmentation [49, 69, 10, 54, 53], etc. In recent years, the scale of vision transformers has grown to be billion-parameters [65, 29]. Consequently, adapting such models with vast scales into downstream tasks presents increasingly complex challenges, particularly in real-world deployment scenarios. Two critical concerns have been widely acknowledged as primary obstacles in implementing large-scale transformers for downstream applications [23, 17, 27]: (1) the training overhead when fine-tuning on downstream tasks, and (2) the inference efficiency after model deployment.
@@
Specifically, first, conventional fine-tuning methods, which necessitate adjusting all parameters of the model (i.e., Full Fine-tuning), suffer from unaffordable consumption of GPU resources and training time given the extensive scales of the foundation models [15, 65, 29]. Researchers have delved into Parameter-Efficient Fine-Tuning (PEFT) [23, 26, 36, 24, 17, 42, 19] algorithms, which generally freeze the pre-trained models and only tune extra small parameters, leading to great reduction of training time and storage overhead. The second issue pertains to inference efficiency, requiring the deployed model to promptly process the input data. The computational complexity of the models significantly influences achieving satisfactory inference throughput. Representative solutions for this matter encompass model compression methods, including model pruning [9, 63, 57], knowledge distillation [21, 1, 51], model quantization [12, 37, 3], etc.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) for Task Adaptation. Transferring large-scale transformers to downstream tasks has been a popular topic in computer vision [26, 28, 36, 67, 68, 8, 17, 59]. PEFT methods either locate a subset of parameters inside the model for fine-tuning [64], or inject human designed modules to the original model structure. Specifically, adapters [23, 48, 28] are a type of MLP module with a bottleneck in the middle. Prompt-tuning [32, 33, 39, 13, 26] insert learnable tokens to model input to generate task-specific outputs. SSF [36] tunes additional scaling and shifting parameters. LoRA [24], AdaptFormer [8], and Consolidator [17] add lightweight modules as bypasses. These modules can be merged with the original backbone for no extra inference cost. In the scope of training-inference efficient task adaptation, PEFT methods retain model inference cost, resulting in extreme difficulty for model deployment. To solve the problem, we conduct adaptive token merging for PEFT. While achieving promising adaptation performance under both low and high compression rates, our method inherits the advantages of PEFT methods.
@@ 3.1 Preliminaries
Then, for each $t^{l}_{i_{\bullet}}$ token, the most similar $t^{l}_{j_{\bullet}}$ token is matched to it via cosine similarity. From $G^{l}_{1}$, $r$ tokens with the most similar connections to tokens in $G^{l}_{2}$ are selected to form token pairs $(t^{l}_{m_{k}},t^{l}_{n_{k}})$, where $t^{l}_{m_{k}}\in G^{l}_{1}$, $t^{l}_{n_{k}}\in G^{l}_{2}$, and $k=1,2,\cdots,r$. Formally,
@@ 3.2 PYRA: Parallel Yielding Re-Activation
Here, we propose a novel Parallel Yielding Re-Activation method (PYRA) to adaptively modulate token features to enhance the perception of data distribution during token merging. Specifically, inside PYRA, the weights for adaptive merging are first yielded in a parallel manner through a pair of lightweight learnable vectors in each ViT block. These generated weights are then applied to modulate tokens to be merged through re-activation. As a result, PYRA enables adaptive calibration of the learned feature distribution with low computational complexity.
@@
Specifically, for layer $l$ with $r$ tokens to be merged, we create two learnable vectors as the modulation weight generator inside the transformer block: $W^{l}_{r}\in\mathbb{R}^{r\times 1}$ and $W^{l}_{D}\in\mathbb{R}^{1\times D}$. To guarantee that $W^{l}_{r}$ and $W^{l}_{D}$ digest the token features from both tokens in a token pair, we first calculate the token information matrix:
@@ 4.1 Experimental Setting
4.2 Performance comparison on Task Adaptation Benchmark
We verify the effectiveness of PYRA for training-inference efficient task adaptation on (1) low compression rate: comparing the performance of different methods under the same sparsity ratio (here we set sparsity ratio=50%); (2) high compression rate: comparing the performance between the competing methods and the smaller-scale model with similar throughput levels. We leverage ViT-L/16 and ViT-B/16 for these inspections. The competing methods include RaP [35], SPViT [30], DiffRate [7], and ToMe [2]. For RaP and SPViT, we first train the backbone with LoRA on downstream tasks, then we prune the models, and lastly we re-train the parameters attached by the pruning method and LoRA. DiffRate and ToMe can be employed with LoRA during fine-tuning, while DiffRate demands ImageNet-21K for searching the optimal compression schedule.
@@ 4.2 Performance comparison on Task Adaptation Benchmark
PYRA on high compression rate. Results on high compression rate are reported in Tab. 3. With comparable throughputs to the smaller-scale model and minimal training parameters, PYRA outperforms all competing methods and achieves the best performance on 35 of 38 dataset metrics. Compared to the smaller-scale model with comparable throughput, PYRA successfully outperforms it. On the compressed ViT-L/16, PYRA even surpasses ViT-B/16 by 0.9%. This shows that PYRA, as the state-of-the-art, effectively bridges the adverse compression gap between compressed large-scale models and small-scale models, as mentioned in Sec. 1. Therefore, as a training-inference efficient method, PYRA is an applicable alternative for acquiring a smaller-scale model (3.2$\times$ speedup) on downstream tasks through efficient training (0.4% training parameters) when no pre-trained small-scale model is available.
@@
PYRA consistently yields better models of different throughputs. To further show the superiority of PYRA for different compression rates, we compress the chosen backbones for a series of speedups and compare the results with the strongest baseline, ToMe [2], under similar values of throughputs. The results are shown in Fig. 2 and Fig. 4. Across all tested compression rates, PYRA consistently outperforms ToMe and flattens the curve of accuracy-throughput for training-inference efficient task adaptation. Therefore, our PYRA is applicable to acquire task-specific smaller-scale models of different throughputs consistently in the absence of pre-trained parameters for smaller-scale models.
@@ 4.4 Further Analysis on Different Designs
We conduct further analysis on ViT-B/16 model pre-trained on ImageNet-21K. The baseline here refers to training LoRA with token merging [2, 7] attached.

### v4

@@ Abstract
Recently, the scale of transformers has grown rapidly, which introduces considerable challenges in terms of training overhead and inference efficiency in the scope of task adaptation. Existing works, namely Parameter-Efficient Fine-Tuning (PEFT) and model compression, have separately investigated the challenges. However, PEFT cannot guarantee the inference efficiency of the original backbone, especially for large-scale models. Model compression requires significant training costs for structure searching and re-training. Consequently, a simple combination of them cannot guarantee accomplishing both training efficiency and inference efficiency with minimal costs. In this paper, we propose a novel Parallel Yielding Re-Activation (PYRA) method for such a challenge of training-inference efficient task adaptation. PYRA first utilizes parallel yielding adaptive weights to comprehensively perceive the data distribution in downstream tasks. A re-activation strategy for token modulation is then applied for tokens to be merged, leading to calibrated token features. Extensive experiments demonstrate that PYRA outperforms all competing methods under both low compression rate and high compression rate, demonstrating its effectiveness and superiority in maintaining both training efficiency and inference efficiency for large-scale foundation models. Our code is available at https://github.com/THU-MIG/PYRA.
@@ 1 Introduction
Vision transformers [21] have made a profound impact across various domains of computer vision, such as image classification [21, 8, 77, 90, 20, 71], object detection [6, 98, 56, 69, 68], image segmentation [70, 94, 13, 75, 96], etc. In recent years, the scale of vision transformers has grown to be billion-parameters [90, 42, 85]. Consequently, adapting such models with vast scales into downstream tasks presents increasingly complex challenges, particularly in real-world deployment scenarios. Two critical concerns have been widely acknowledged as primary obstacles in implementing large-scale transformers for downstream applications [33, 26, 37, 67]: (1) the training overhead when fine-tuning on downstream tasks, and (2) the inference efficiency after model deployment.
@@
Specifically, first, conventional fine-tuning methods, which necessitate adjusting all parameters of the model (i.e., Full Fine-tuning), suffer from unaffordable consumption of GPU resources and training time given the extensive scales of the foundation models [21, 90, 42]. Researchers have delved into Parameter-Efficient Fine-Tuning (PEFT) [33, 36, 51, 34, 26, 28] algorithms, which generally freeze the pre-trained models and only tune extra small parameters, leading to great reduction of training time and storage overhead. The second issue pertains to inference efficiency, requiring the deployed model to promptly process the input data. The computational complexity of the models significantly influences achieving satisfactory inference throughput. Representative solutions for this matter encompass model compression methods, including model pruning [12, 88, 79, 9], knowledge distillation [31, 1, 73, 80, 81], model quantization [17, 53, 5], etc.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) for Task Adaptation. Transferring large-scale transformers to downstream tasks has been a popular topic in computer vision [36, 38, 51, 92, 93, 11, 26, 84]. PEFT methods either locate a subset of parameters inside the model for fine-tuning [89], or inject human designed modules to the original model structure. Specifically, adapters [33, 66, 38, 58] are a type of MLP module with a bottleneck in the middle. Prompt-tuning [47, 48, 55, 19, 36, 27] insert learnable tokens to model input to generate task-specific outputs. SSF [51] tunes additional scaling and shifting parameters. LoRA [34], AdaptFormer [11], and Consolidator [26] add lightweight modules as bypasses. These modules can be merged with the original backbone for no extra inference cost. In the scope of training-inference efficient task adaptation, PEFT methods retain model inference cost, resulting in extreme difficulty for model deployment. To solve the problem, we conduct adaptive token merging for PEFT. While achieving promising adaptation performance under both low and high compression rates, our method inherits the advantages of PEFT methods.
@@ 3.1 Preliminaries
Then, for each $t^{l}_{i_{\bullet}}$ token, the most similar $t^{l}_{j_{\bullet}}$ token is matched to it via cosine similarity. From $G^{l}_{1}$, $r$ tokens with the most similar connections to tokens in $G^{l}_{2}$ are selected to form token pairs $(t^{l}_{m_{k}},t^{l}_{n_{k}})$, where $t^{l}_{m_{k}}\in G^{l}_{1}$, $t^{l}_{n_{k}}\in G^{l}_{2}$, and $k=1,2,\cdots,r$. Note that $\{m_{k}\}$ is a re-indexed subset of $\{i_{\bullet}\}$. Formally,
@@ 3.2 PYRA: Parallel Yielding Re-Activation
Here, we propose Parallel Yielding Re-Activation (PYRA) to adaptively modulate token features to enhance the perception of data distribution during token merging. Specifically, inside PYRA, the weights for adaptive merging are first yielded in a parallel manner through a pair of lightweight learnable vectors in each ViT block. These generated weights are then applied to modulate tokens to be merged through re-activation. As a result, PYRA enables adaptive calibration of the learned feature distribution with low computational complexity.
@@
Specifically, for layer $l$ with $r$ tokens to be merged, we create two learnable vectors as the modulation weight generator inside the transformer block: $W^{l}_{r}\in\mathbb{R}^{r\times 1}$ and $W^{l}_{D}\in\mathbb{R}^{1\times D}$. To guarantee that $W^{l}_{r}$ and $W^{l}_{D}$ digest the token features from both tokens in a token pair, we first normalize the sum of token pairs to construct the token information matrix:
@@ 4.1 Experimental Setting
4.2 Performance comparison on Task Adaptation Benchmark
We verify the effectiveness of PYRA for training-inference efficient task adaptation on (1) low compression rate: comparing the performance of different methods under the same sparsity ratio (here we set sparsity ratio=50%); (2) high compression rate: comparing the performance between the competing methods and the smaller-scale model with similar throughput levels. We leverage ViT-L/16 and ViT-B/16 for these inspections. The competing methods include RaP [50], SPViT [43], DiffRate [10], and ToMe [3]. Comparisons to more baselines are in the supplementary materials. For RaP and SPViT, we first train the backbone with LoRA on downstream tasks, then we prune the models, and lastly we re-train the parameters attached by the pruning method and LoRA. DiffRate and ToMe can be employed with LoRA during fine-tuning, while DiffRate demands ImageNet-21K for searching the optimal compression schedule.
@@ 4.2 Performance comparison on Task Adaptation Benchmark
PYRA on high compression rate. Results are reported in Tab. 3. With comparable throughputs to the smaller-scale model and minimal training parameters, PYRA outperforms all competing methods and achieves the best performance on 35 of 38 dataset metrics. Compared to the smaller-scale model with comparable throughput, PYRA successfully outperforms it. On the compressed ViT-L/16, PYRA even surpasses ViT-B/16 by 0.9%. This shows that PYRA, as the state-of-the-art, effectively bridges the adverse compression gap between compressed large-scale models and small-scale models, as mentioned in Sec. 1. Therefore, PYRA is an applicable alternative for acquiring a smaller-scale model (3.2$\times$ speedup) on downstream tasks through efficient training (0.4% training parameters) when no pre-trained small-scale model is available.
@@
PYRA consistently yields better models of different throughputs. To further show the superiority of PYRA for different compression rates, we compress the chosen backbones for a series of speedups and compare the results with the strongest baseline, ToMe [3], under similar values of throughputs. As shown in Figs. 2 and 4, PYRA consistently outperforms ToMe and flattens the accuracy-throughput curve. Therefore, our PYRA is applicable to acquire task-specific smaller-scale models of different throughputs consistently in the absence of pre-trained parameters for smaller-scale models.
@@ 4.4 Further Analysis on Different Designs
We conduct further analysis on ViT-B/16 model pre-trained on ImageNet-21K. The baseline here refers to training LoRA with token merging [3, 10] attached. More analysis experiments can be found in the supplementary materials.
@@
5 Conclusion
Limitations. We have not yet validated the effectiveness of PYRA in object detection and image segmentation. We plan to extend these tasks in the future. Meanwhile, it is also applicable to attach pruning methods after training the vision transformers with PYRA to eliminate additional parameters attached by PYRA. We leave that topic to future studies.
Acknowledgements
This work was supported by National Science and Technology Major 2022ZD0119401, National Natural Science Foundation of China (Nos. 62271281, 61925107, 62021002). It is also sponsored by CAAI-CANN Open Fund, developed on OpenI Community.
Supplementary Material
Table 9: The statistics of the VTAB-1k [91] benchmark.
Dataset Description Classes Train size Val size Test size
CIFAR-100 [45] Natural 100 800/1000 200 10000
Caltech101 [23] 102 6084
DTD [16] 47 1880
Flowers102 [63] 102 6149
Pets [64] 37 3669
SVHN [62] 10 26032
Sun397 [83] 397 21750
Patch Camelyon [74] Specialized 2 800/1000 200 32768
EuroSAT [30] 10 5400
Resisc45 [14] 45 6300
Retinopathy [25] 5 42670
Clevr/count [40] Structured 8 800/1000 200 15000
Clevr/distance [40] 6 15000
DMLab [2] 6 22735
KITTI/distance [24] 4 711
dSprites/location [61] 16 73728
dSprites/orientation [61] 16 73728
SmallNORB/azimuth [46] 18 12150
SmallNORB/elevation [46] 9 12150
A Details for the Evaluation Datasets
We show the detailed statistics of the Visual Task Adaptation Benchmark (VTAB-1k) [91] in Tab. 9. Introduced in [91], the VTAB-1k benchmark contains 19 tasks from diverse domains: (1) Natural images that are captured by standard cameras in real-world scenarios; (2) Specialized images that are captured by professional equipment, e.g., remote sensing and medical cameras; (3) Structured images that are synthesized from simulated environments. VTAB-1k contains a variety of tasks such as object counting and depth estimation apart from the standard image classification. For each task in VTAB-1k, the images are divided into the training set (800 images), the validation set (200 images), and the test set (the original set). All models are fine-tuned on the train+val set, i.e., 1000 images. The validation set is used for hyperparameter tuning.
Table 10: PEFT hyperparameters for each tested backbone.
ViT-B/16 ViT-L/16 ViT-L/16 (MAE) DeiT-B/16
LoRA $h$ 8 12 12 8
Table 11: Training hyperparameters for each tested backbone.
ViT-B/16 ViT-L/16 ViT-L/16 (MAE) DeiT-B/16
optimizer AdamW AdamW AdamW AdamW
warmup epochs 10 10 10 10
epochs 100 100 100 100
batch size 64 32 32 128
lr (PEFT) 1e-3 1e-3 1e-3 1e-3
wd 1e-4 1e-4 1e-4 1e-4
Table 12: The learning rate choices for $W_{D}$ and $W_{r}$ in each tested backbone.
lr choices for $W_{D}$ and $W_{r}$
ViT-B/16 [1e-5, 3e-5, 1e-4, 3e-4, 1e-3, 3e-3]
ViT-L/16 [1e-5, 3e-5, 1e-4, 3e-4, 1e-3, 3e-3]
ViT-L/16 (MAE) [1e-6, 3e-6, 1e-5, 3e-5]
DeiT-B/16 [1e-5, 3e-5, 1e-4, 3e-4, 1e-3, 3e-3]
Table 13: The merging schedules for the high compression rate benchmark.
$r$ schedule ([$r^{1}$, $r^{2}$$\cdots$, $r^{L}$])
ViT-B/16 [40,34,30,24,18,14,10,8,4,4,3,3]
ViT-L/16 [20,19,18,17,15,13,13,12,10,9,8,6,6,4,4,4,3,3,2,2,1,1,1,1]
ViT-L/16 (MAE) [20,19,18,17,15,13,13,12,10,9,8,6,6,4,4,4,3,3,2,2,1,1,1,1]
DeiT-B/16 [40,34,30,24,18,14,10,8,4,4,3,3]
B More Implementation Details
We present details of implementing PYRA for training-inference efficient task adaptation.
PEFT Hyperparameters. For training-inference efficient task adaptation, we implement LoRA [34] with hidden layer dimension $h$ as the PEFT module regarding its simplicity and mergeability. We present the PEFT hyperparameters in Tab. 10. Note that $h$ choices in Tab. 10 lead to approximately the same percentage of training parameters compared to the backbone for all vision transformers. We use the same $h$ value for each backbone under all competing methods and our PYRA.
Training Hyperparameters. For choices of training hyperparameters, we mainly follow [93, 26] and only conduct minor adjustments on the training batch size for efficient training on our devices. The hyperparameters are listed in Tab. 11. To prevent the token modulation from overly altering tokens, thereby leading to training instability, we adopt different learning rates for the modulation weight generators $W_{D}$ and $W_{r}$ introduced in PYRA. Specifically, for each tested model, we tune for the optimal learning rate for $W_{D}$ and $W_{r}$ on each task within the corresponding range as listed in Tab. 12.
Changing Compression Rate for PYRA. To yield compressed smaller-scale models with different throughput values, we change the compression rate of PYRA by changing the $r$ value, i.e., merged token number, for each layer.
(1) For the low compression rate benchmark, we simply implement a constant merging schedule which simply merges the same number of tokens for each layer. Specifically, we merge 16 tokens in each layer for ViT-B/16 and DeiT-B/16, and merge 8 tokens in each layer for ViT-L/16 and ViT-L/16 (MAE).
(2) For the high compression rate benchmark, we implement a decreasing schedule that consistently outperforms other schedules on the pre-trained dataset [3]. To achieve the target throughput value demanded by the high compression rate benchmark, we solve an approximate problem and apply minor adjustments to the solutions based on actually tested throughput values. Formally, for a vision transformer with $L$ layers, total token number $T$ (here we omit the [CLS] token and the distillation token), and $F$ computational FLOPs, we compress it to a smaller-scale model with $f$ FLOPs by merging $r^{l}$ tokens in layer $l$. We denote the number of remaining tokens in the last layer as $t$ (usually $t<5$). We calculate the approximate solution $\hat{r}^{l}$ for the merging schedule as:
$\hat{r}^{l}=\lfloor(g(l-1)-g(l))(T-t)\rfloor,\text{where}\ g(x)=(1-\frac{x}{L})^{\frac{F}{f}-1}.$ (8)
After acquiring the approximate solutions $\hat{r}^{l}$, we apply minor adjustments based on the actually tested throughput values to yield the demanded speedup. The final adjusted $r^{l}$ yields a smaller-scale compressed model with similar throughput values as the smaller-scale backbone, as listed in Tab. 3 of the article. The actually applied $r^{l}$ values for the high compression rate benchmark are listed in Tab. 13. Note that Eq. 8 can also yield the adopted constant merging schedule for the low compression rate benchmark.
For benchmark results in Sec. 4.2 of the article, we adopt the same merging schedules for PYRA and ToMe [3] to guarantee a fair comparison.
Table 14: Comparison of different token pruning method choices on the ViT-B/16 with LoRA modules attached as the PEFT choice.
Compression Rate Method # params Throughput Average
Low EViT[52] 0.34% 732 73.09
ToFu[41] 0.34% 748 73.31
ATS[22] 0.34% 727 71.16
LTMP[4] 0.34% 724 52.70
ToMe[3] 0.34% 753 74.10
PYRA 0.35% 745 74.69
High EViT[52] 0.34% 1200 65.79
ToFu[41] 0.34% 1370 70.39
ATS[22] 0.34% 1183 61.12
LTMP[4] 0.34% 1265 48.96
ToMe[3] 0.34% 1381 70.43
PYRA 0.35% 1365 72.06
C Comparing to More Baseline Methods
Apart from the baseline methods chosen in the main article, we applied several more token pruning methods to demonstrate that our choice of ToMe[3] is a strong baseline in the setting of training-inference efficient task adaptation. Specifically, we have applied EViT[52], ToFu[41], ATS[22], and LTMP[4] with LoRA attached as the learnable PEFT module. As shown in Tab. 14, all listed methods underperform the strongest baseline, ToMe[3], well demonstrating that analyzing and evaluating ToMe in training-inference efficient task adaptation is important and appropriate. Note that although [41, 4] are follow-up works of ToMe, they involve heavy token pruning that requires extensive training to restore performance, hence underperforming ToMe in PEFT training.
D Comparing to Parameter-Efficient Model Distillation
Table 15: Comparison to PE-Dist under the high compression rate.
Backbone Method # params Throughput Natural Specialized Structured Average
ViT-S/16 PEFT 0.34% 1350 76.29 83.56 55.71 71.85
PE-Dist 0.34% 1350 76.29 83.84 52.85 70.99
ViT-B/16 PYRA 0.35% 1365 73.91 82.60 59.66 72.06
ViT-B/16 PEFT 0.34% 425 79.45 84.43 60.39 74.76
PE-Dist 0.34% 425 80.30 84.24 59.26 74.60
ViT-L/16 PYRA 0.40% 427 80.43 85.17 61.39 75.66
Model distillation [31] is an alternative approach for acquiring a smaller-scale model of the desired size. Existing approaches to model distillation [31, 1, 73] require training all parameters in the smaller-scale model with a combined loss:
$\mathcal{L}=\mathcal{L}_{\text{task}}+\mathcal{L}_{\text{dist}},$ (9)
where $\mathcal{L}_{\text{task}}$ represents the task-specific loss, and $\mathcal{L}_{\text{dist}}$ denotes the distillation loss, typically measured by the KL-divergence.
In the setting of training-inference efficient task adaptation, however, training all parameters in the smaller-scale model is unacceptable. Therefore, we implement a baseline as parameter-efficient model distillation (PE-Dist), in which we attach PEFT modules to existing pre-trained smaller-scale backbones, and only train the PEFT modules during model distillation. We conduct experiments on the ViT backbones [21], i.e., ViT-B/16 and ViT-L/16, by distilling them to the smaller-scale backbone, i.e., ViT-S/16 and ViT-B/16, correspondingly. To yield the distillation source, we first train ViT-B/16 and ViT-L/16 by PEFT. As a fair comparison, we choose LoRA [34] as the PEFT method and compare the results for PE-Dist with our PYRA under the high compression rate. As shown in Tab. 15, parameter-efficient model distillation consistently underperforms directly fine-tuning the smaller-scale model for both backbones, indicating that the adverse compression phenomenon still exists when applying model distillation for training-inference efficient task adaptation. In contrast to model distillation, our PYRA consistently eliminates the adverse compression gap and outperforms model distillation by over >1% on both backbones as well. This indicates that PYRA represents an effective alternative to model distillation for acquiring smaller-scale inference-efficient models under the constraints of parameter-efficient training.
E More Analysis Experiments.
Table 16: Comparison of different pipelines on the ViT-B/16.
Compression Rate Pipeline Method # params Throughput Average
Low One-Stage ToMe 0.34% 753 74.10
Two-Stage 0.34% 753 72.73
High One-Stage ToMe 0.34% 1381 70.43
Two-Stage 0.34% 1381 69.25
Is a Two-Stage Schedule Better? As discussed in Sec. 3.1 in the article, we employ a one-stage pipeline in which LoRA modules are trained while token merging is attached to the vision transformer. In fact, a more straightforward approach is to implement in a two-stage pipeline: we train the LoRA modules without token merging first, and then employ token merging in the second stage to reduce the redundant tokens. To compare the performance of both pipelines, we apply the ViT-B/16 model pre-trained on ImageNet-21K [18] with ToMe attached as the compression method. As shown in Tab. 16, under both low and high compression rates, the one-stage pipeline described in the main article outperforms the two-stage pipeline, well demonstrating that our pipeline choice is appropriate.